K. Ison (beginning capital, $42,000) and I. McCoy (beginning capital $67,200) are partners. During 2021, the partnership reported profit of $50,400, and Ison made drawings of $12,600 while McCoy made drawings of $16,800.

Assume the partnership agreement calls for profit to be divided 55% to Ison and 45% to McCoy. Prepare the journal entry to close the Income Summary account. (Credit account titles are automatically indented when the amount is entered. Do not indent manually.)

Account Titles and Explanation

Debit

Credit

    I. McCoy, Capital    Income Summary    K. Ison, Capital    

    K. Ison, Capital    Income Summary    I. McCoy, Capital    

    I. McCoy, Capital    K. Ison, Capital    Income Summary    

(To close Income Summary.)
Assume the partnership agreement calls for profit to be divided with a salary of $25,200 to K. Ison and $16,800 to I. McCoy, with the remainder divided 55% to Ison and 45% to McCoy. Prepare the journal entry to close the Income Summary account. (Credit account titles are automatically indented when the amount is entered. Do not indent manually.)

Account Titles and Explanation

Debit

Credit

    I. McCoy, Capital    K. Ison, Capital    Income Summary    

    I. McCoy, Capital    K. Ison, Capital    Income Summary    

    Income Summary    K. Ison, Capital    I. McCoy, Capital    

(To close Income Summary.)
Assume the partnership agreement calls for profit to be divided with a salary of $33,600 to K. Ison and $25,200 to I. McCoy, interest of 10% on beginning capital, and the remainder divided 50%–50%. Prepare the journal entry to close the Income Summary account. (Credit account titles are automatically indented when the amount is entered. Do not indent manually.)

Account Titles and Explanation

Debit

Credit

    K. Ison, Capital    I. McCoy, Capital    Income Summary    

    K. Ison, Capital    I. McCoy, Capital    Income Summary    

    I. McCoy, Capital    Income Summary    K. Ison, Capital    

(To close Income Summary.)
Answer #1

1)

Date Account title Debit credit
Income summary 50400
K. Ison, Capital      (50400*55%) 27720
I. McCoy, Capital (50400*45%) 22680

2)

K. Ison, Capital I. McCoy, Capital Total
Salary 25200 16800 42000
Remaining profit of 50400-42000= 8400 in 55% and 45% 8400*55%= 4620 3780 8400
Total 29820 20580 50400
Date Account title Debit credit
Income summary 50400
K. Ison, Capital      29820
I. McCoy, Capital 20580

3)

K. Ison, Capital I. McCoy, Capital Total
Salary 33600 25200 58800
Interest on beginning capital 42000*10%=4200 67200*10%=6720 10920
Remaining profit/(loss) of 50400--58800-10920= -19320 in 50% and 50% -19320*50%=-9660 -9660 -19320
Total 28140 22260 50400
Date Account title Debit credit
Income summary 50400
K. Ison, Capital      28140
I. McCoy, Capital 22260